# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

### v2

@@ Abstract
The bounds (8)-(9) may require solving linear systems involving very large covariance matrices. For multipartite systems, it is often possible to split these bounds into $N$ bounds that involve smaller matrices (see End Matter).
Related work.— Eq. (3) is closely related to the Donsker-Varadhan representation from large deviations theory [23], previously used to estimate information-theoretic quantities in machine learning [24] and thermodynamic inference [25, 26, 27]. However, these works did not consider the problem of thermodynamic inference in high-dimensional systems using multivariate observables. Other approaches to EP inference have employed compression algorithms [28, 29, 30] and deep learning of probability flows [31].
Our Pythagorean theorem (5) is related to previous decompositions of EP in interacting systems [32] and systems with nonconservative forces [33, 34, 15]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [35, 36, 37, 38], sometimes called “maximum caliber” (MaxCal) [39]. Related techniques have been used to infer models from which the entropy flow can be estimated, though without ensuring a lower bound [40].
Our approach differs in several other ways from earlier work on MaxEnt and MaxCal. First, we minimize the KL divergence relative to an unknown prior distribution $\tilde{p}$, from which we typically only have some samples. This makes standard approximation methods for large-scale MaxEnt problems (such as mean-field and Bethe approximations [41, 42, 6]) not directly applicable. Second, we care not only about the parameters ${\bm{\theta}}$, as is typical of MaxEnt “inverse problems”, but also about bounding EP by the quantity ${\Sigma}_{\bm{g}}$. Both of these issues are resolved by the dual formulation (3).
Lynn et al. [43, 44, 12] proposed a way to decompose and bound EP using a different information-theoretic optimization. Although originally focused on local EP in multipartite stationarity systems, it can be generalized to arbitrary observables $\bm{g}$ and non-multipartite systems as
@@
${\Sigma}_{\bm{g}}^{\textrm{L}}$ $=\min_{q,\tilde{q}}D(q\|\tilde{q})\;\>\text{subject to}\;\>\langle\bm{g}\rangle_{q}=\langle\bm{g}\rangle_{p}.$ (10)
Here we use assumption of stationarity and no odd variables, so that $\tilde{q}(\vec{{\bm{x}}})=q(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$. Like our bound ${\Sigma}_{\bm{g}}$, ${\Sigma}_{\bm{g}}^{\textrm{L}}$ can be used to generate hierarchical decompositions of EP.
@@
Although Eq. (10) resembles Eq. (2), it differs in that it simultaneously optimizes both arguments of the KL divergence, $q$ and $\tilde{q}$. As we discuss in the End Matter, the optimal distribution $q^{*}$ in Eq. (10) is not in an exponential family and in general it lacks full support. For this reason, Eq. (10) does not have a tractable dual expression (i.e., analogous to our dual (3)), and there is no straightforward way to scale ${\Sigma}_{\bm{g}}^{\textrm{L}}$ to large systems.
@@
Examples: Nonequilibrium spin model and Neuropixels dataset.— We illustrate our approach using two examples: a nonequilibrium kinetic Ising model [45, 46, 6], and in vivo spike data from the Neuropixels Visual Behavior repository [47]. In both systems, we focus on a single time step and binary variables $x_{i,t}\in\{-1,+1\}$, $i\in\{1,\dots,N\}$ with time $t\in\{0,1\}$, and assume stationarity.
@@
To calculate ${\Sigma}_{\bm{g}}$, we optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [48]. We perform early-stopping using held-out validation data, which avoids overfitting even in the far-from-equilibrium regime where many reverse transitions are not sampled. Reported EP estimates are computed on held-out test data. Details of the data generation, optimization, and analysis are found in the SM [16] and our code repository [49].
@@
Here we introduce the spin-flip operator, $(\bm{x}^{[i]})_{i}=-x_{i}$ and $(\bm{x}^{[i]})_{j}=x_{j}$ for $j\neq i$. The flip probability for spin $i$ is
@@
$g_{ij}(\vec{{\bm{x}}})=(x_{i,1}-x_{i,0})x_{j,0}\qquad\text{for all}\quad i,j\,,$ (12)
with conjugate parameters $\theta_{ij}$. Because the system has multipartite dynamics and observables, Eq. (3) can be decomposed into $N$ independent problems, each involving only those transitions where spin $i$ changes, improving computational performance for large systems.
@@
As discussed near Eq. (6), ${\Sigma}_{\bm{g}}\approx{\Sigma}$ implies that the trajectory EP $\sigma(\vec{{\bm{x}}})$ can be closely approximated by a linear combination of observables $g_{ij}$. As shown in the SM, our optimal parameters ${\bm{\theta}}^{*}$ can be used to infer the asymmetry of the coupling constants, $\theta_{ij}^{*}-\theta_{ji}^{*}\approx\beta(w_{ij}-w_{ji})$. This relation is verified in Fig. 1(b).
In the SM, we also consider a different family of coupling matrices, for which this model exhibits a nonequilibrium phase transition [6]. We show that our approach accurately infers EP in this regime, a difficult task for existing approximations [53].
@@
Next, as an example application to biological data, we estimate EP in the Visual Behavior Neuropixels dataset [47], containing spike-train recordings from 81 mice across 103 sessions. EP is here interpreted as a statistical measure of time-irreversibility, not energetic dissipation [13]. The dataset includes spiking activity from multiple brain regions, including visual cortical areas (VISp, VISl, VISal, VISrl, VISam and VISpm) and subcortical structures. We analyze data from visual areas during three conditions: active behavior (visual change detection task), passive replay (same stimuli but without task engagement), and Gabor (receptive field characterization with Gabor stimuli and full-field flashes). We discretize spike trains into temporal bins of length $\Delta t=10$ ms and verify that most bins contain no more than one spike. We note that neurons can update in a parallel (non-multipartite) manner.
@@
$g_{ij}(\vec{{\bm{x}}})=x_{i,1}\,x_{j,0}-x_{i,0}\,x_{j,1}\qquad\text{for all}\quad i<j\,,$ (13)
@@
We calculate EP estimates across different conditions and system sizes $N$. For each combination of size, condition, and recording session, we randomly select 10 subsets of neurons and estimate ${\Sigma}_{\bm{g}}$ for each subset. For improved comparison, we normalize EP in each condition by the expected number of spikes per bin, $R=\sum_{i}\left(1+\langle x_{i,0}\rangle_{\rm st}\right)/2$.
Fig. 2(a) shows that EP grows superlinearly with size (note that $R\propto N$). It also shows that the active condition is associated with the largest normalized EP. Fig. 2(b) illustrates the matrix of inferred parameters $\theta_{ij}^{*}$ for 200 neurons from an active trial, grouped by visual brain area. These parameters can be understood as the coupling coefficients that best describe temporal irreversibility of the in vivo neural dynamics. Notably, they reveal a clustered organization that is aligned with anatomical brain regions.

### v3

@@ Abstract
The bounds (11)-(12) may require solving linear systems involving very large covariance matrices. For multipartite observables, it is often possible to split these bounds into sub-problems that involve smaller matrices (see End Matter).
[…]
where $v_{\bm{\phi}}({{\bm{x}}})$ is a nonlinear function parameterized by $\bm{\phi}\in\mathbb{R}^{k}$ (e.g., output of a neural network)  [47]. ${\Sigma}_{\bm{g}}$ in Eq. (3) is a special case of $\Sigma_{\bm{g}}^{\text{B}}$ where $v_{\bm{\phi}}({{\bm{x}}})$ contains all linear combinations of $\bm{g}({{\bm{x}}})$. The neural-network-based bound (13) has both practical and conceptual differences with respect to Eq. (3). For instance, it generally involves a difficult non-convex optimization problem, it does not provide a Pythagorean decomposition of EP, and it cannot be interpreted in terms of large deviations. Exploration of $\Sigma_{\bm{g}}^{\text{B}}$ for thermodynamic inference is left as an interesting future direction.
A related variational expression for thermodynamic inference was developed in Kim et al. [39] and Otsubo et al. [40]. In our notation, it is written as
@@
$\Sigma_{\bm{g}}^{\text{KO}}\vcentcolon=\max_{{\bm{\theta}}}\left({\bm{\theta}}^{\top}\langle\bm{g}\rangle_{p}-\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}+1\right)\,.$ (14)
It is possible to consider a neural-network-based version of $\Sigma_{\bm{g}}^{\text{KO}}$ by replacing ${\bm{\theta}}^{\top}\bm{g}({{\bm{x}}})$ with a parameterized function $v_{\bm{\phi}}({{\bm{x}}})$, as in Eq. (13), as explored in Refs. [39, 49].
Note that ${\Sigma}_{\bm{g}}\geq\Sigma_{\bm{g}}^{\text{KO}}$, since $\ln x\geq-x+1$ [47, 50], therefore ${\Sigma}_{\bm{g}}$ always provides a tighter bound on EP than $\Sigma_{\bm{g}}^{\text{KO}}$. The two bounds become equivalent in the limit $\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\to 1$, such as the short-time limit $T\to 0$ with antisymmetric observables when $\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}=1+O(T)$ [40]. In general, however, the two bounds give different results, and in the End Matter, we provide a simple example where ${\Sigma}_{\bm{g}}$ is arbitrarily better than $\Sigma_{\bm{g}}^{\text{KO}}$. We note that outside the short-time limit, $\Sigma_{\bm{g}}^{\text{KO}}$ does not have a straightforward interpretation in terms of maximum likelihood inference or large-deviations statistics.
Lynn et al. [51, 52, 32] proposed a way to decompose and bound steady-state EP using a different information-theoretic optimization. Although originally focused on local EP in systems with multipartite dynamics, it can be generalized to arbitrary observables $\bm{g}$ and non-multipartite systems as
[…]
@@
Although Eq. (15) resembles Eq. (2), it differs in that it simultaneously optimizes both arguments of the KL divergence, $q$ and $\tilde{q}$. As discussed in the End Matter, the optimal distribution in Eq. (15) is not in an exponential family, and in general it lacks full support. For this reason, Eq. (15) does not have a tractable dual expression analogous to our dual (3), and there is no straightforward way to scale ${\Sigma}_{\bm{g}}^{\textrm{L}}$ to large systems.
Other work has explored objective (15) in combination with other types of constraints, e.g., on waiting times [19, 14] and hidden Markovian structure [23, 13]. However, the resulting optimization typically does not scale to high-dimensional or even moderately-sized systems, in part because the considered constraints lead to non-convex optimization problems.
Our Pythagorean theorem (6) is related to previous decompositions of EP in interacting systems [53] and systems with nonconservative forces [54, 55, 35]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [56, 57, 58, 59], sometimes called “maximum caliber” (MaxCal) [60]. Related techniques have been used to infer models from which the entropy flow can be estimated, though without ensuring a lower bound on EP [61].
Our approach differs in several other ways from earlier work on MaxEnt and MaxCal. First, we minimize the KL divergence relative to an unknown prior distribution $\tilde{p}$, from which we typically only have samples. This makes standard approximation methods for large-scale MaxEnt problems, such as mean-field and Bethe approximations [62, 63, 26], not directly applicable. Second, we care not only about the parameters $\bm{\theta}$}, as is typical of MaxEnt “inverse problems”, but also about bounding EP by the quantity ${\Sigma}_{\bm{g}}$. Both of these issues are resolved by the dual formulation (3).
Other notable approaches to EP inference have employed compression algorithms [64, 65, 66] and deep learning of probability flows [67].
@@
Examples: Nonequilibrium spin model and Neuropixels dataset.— We illustrate our method on two examples: a nonequilibrium kinetic Ising model [68, 69, 26] and in vivo spike data from the Neuropixels Visual Behavior repository [70]. For both examples, we consider a single discrete time step and binary variables, ${{\bm{x}}}=(\bm{x}_{0},\bm{x}_{1})$ where $x_{i,t}\in\{-1,+1\}$ for $i\in\{1,\dots,N\}$. We emphasize that our approach also works for continuous systems. In the SM [37], we illustrate it on a linear Langevin system.
@@
To calculate ${\Sigma}_{\bm{g}}$, we sample from steady state using Monte Carlo. We then optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [71]. We perform early-stopping using held-out validation data, which avoids overfitting even in the far-from-equilibrium regime where many reverse transitions are not sampled. Reported EP estimates are computed on held-out test data. Details of the data generation, optimization, and analysis can be found in the SM [37] and our code repository [72].
@@
where we introduced the spin-flip operator: $(\bm{x}^{[i]})_{i}=-x_{i}$ and $(\bm{x}^{[i]})_{j}=x_{j}$ for $j\neq i$. The flip probability for spin $i$ is
@@
$g_{ij}({{\bm{x}}})=(x_{i,1}-x_{i,0})x_{j,0}\qquad\text{for all}\quad i,j\,$, (17)
with conjugate parameters $\theta_{ij}$. Because the system has multipartite observables, Eq. (3) can be decomposed into $N$ independent problems, each involving only those transitions where spin $i$ changes, improving computational performance for large systems (see SM).
@@
As discussed above, ${\Sigma}_{\bm{g}}\approx{\Sigma}$ implies that the trajectory EP $\sigma({{\bm{x}}})$ is closely approximated by $\sigma_{{\bm{\theta}}^{*}}$ from Eq. (8). As shown in the SM, this allows us to use the optimal parameters $\bm{\theta}$$^{*}$ to infer the asymmetry of the coupling constants, $\theta_{ij}^{*}-\theta_{ji}^{*}\approx\beta(w_{ij}-w_{ji})$. This relation is verified in Fig. 1(b).
In the SM, we also consider other coupling matrices for which the model exhibits a nonequilibrium phase transition [26]. We show that our method accurately infers EP in this regime, a difficult task for existing approximations [76].
@@
Next, as an example application to biological data, we estimate EP in the Visual Behavior Neuropixels dataset [70], which contains spike-train recordings from 81 mice in 103 sessions. In this example, EP is understood as a statistical measure of temporal irreversibility, rather than energetic dissipation [33]. The dataset includes spiking activity from multiple brain regions, including visual cortical areas (VISp, VISl, VISal, VISrl, VISam and VISpm) and subcortical structures.
We analyze data from visual areas during three conditions: active behavior (visual change detection task), passive replay (same stimuli but without task engagement), and Gabor (receptive field characterization with Gabor stimuli and full-field flashes). We discretize spike trains into temporal bins of length $10$ ms and verify that most bins contain no more than one spike. In this dataset, neurons can update in a parallel (non-multipartite) manner.
@@
$g_{ij}({{\bm{x}}})=x_{i,1}\,x_{j,0}-x_{i,0}\,x_{j,1}\qquad\text{for all}\quad i<j\,$, (18)
@@
We estimate EP across different conditions, recording sessions, and system sizes $N$. We randomly select 10 subsets of neurons and estimate ${\Sigma}_{\bm{g}}$ for each subset. For improved comparison, we normalize EP in each condition by the expected number of spikes per bin, $R=\sum_{i}\left(1+\langle x_{i,0}\rangle_{\rm st}\right)/2$.
Fig. 2(a) shows that EP grows superlinearly with size (note that $R\propto N$). It also shows that the active condition is associated with the largest normalized EP. Fig. 2(b) illustrates the matrix of inferred parameters $\theta_{ij}^{*}$ for 200 neurons from an active trial, grouped by visual brain area. This matrix specifies a network of functional connectivity [77], showing how interactions between individual neurons contribute to temporal irreversibility of in vivo brain dynamics. This network reveals a clustered organization aligned with anatomical brain regions.